OpenStudy (anonymous):

Who can help me find area between curves? 7x+y^2=8 x=y i dont know how to approach this one.

OpenStudy (amistre64):

the area between 2 objects can be found by subtracting the area of one from the other

OpenStudy (amistre64):

we need to know, or determine, where these "curves" are the same; where they cross at

OpenStudy (amistre64):

it might be better to read if we use "x" is the dependant variable

OpenStudy (amistre64):

i cant really see away to help you thru this if you aint gonna participate :/

OpenStudy (anonymous):

sorry, i was waiting and went to eat, now im back, and i know that i need to subtract area from of 1 from another, but how do i know the enclosed region?

OpenStudy (amistre64):

the enclosed region is where the curves meet and "enclose" the area 7x + y^2 = 8 x =y ; solve this system of equations, perhaps by multiplying the bottom by -7 7x + y^2 = 8 -7x = -7y --------------- y^2 = 8-7y ; +7y and -8 +7y -8 -8 +7y ------------------ y^2 +7y -8 = 0 factor and solve for y

OpenStudy (anonymous):

thanks, y =-1 ; y=8

OpenStudy (anonymous):

oppps, missed signs, -8,1

OpenStudy (amistre64):

good :) so we integrate each equation from -8 to 1; and subtract one from the other. if you end up with a negative it just means you subtracted in the wrong order .... take the absolute value of the result and youll be fine

OpenStudy (amistre64):

since x=y is already in good form, lets modify the other one to match it up 7x + y^2 = 8 x = (8 - y^2)/7

OpenStudy (amistre64):

now that everything is speaking in terms of "y"; the integration does easier